OpenStudy (anonymous):

Find m of angle a if a = 34, b = 28 and c = 15. (Points : 3)

OpenStudy (anonymous):

as in a triangle? need law of cosines for this one

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you can start with \[a^2=b^2+c^2-2bc\cos(A)\] and solve for \(\cos(A)\) or go right to \[\cos(A)=\frac{b^2+c^2-a^2}{2bc}\]

OpenStudy (anonymous):

so for angle A you need \[\cos^{-1}(\frac{b^2+c^2-a^2}{2bc})\]

OpenStudy (anonymous):

calculator exercise now

OpenStudy (anonymous):

ok ummm can u show me

OpenStudy (anonymous):

replace a by 34, b by 28, c by 15

OpenStudy (anonymous):

ok what is the cos thing?

OpenStudy (anonymous):

inverse cosine or arccosine here is what i got from wolfram http://www.wolframalpha.com/input/?i=arccos%28%2824%5E2%2B25%5E2-34%5E2%29%2F%282*15*24%29%29

OpenStudy (anonymous):

i am not sure what you are asking. this problem is solved using the law of cosines. so in order to do it you need to know about cosines

OpenStudy (anonymous):

ok so the answer is 2.329?

OpenStudy (anonymous):

no, check the link i sent

OpenStudy (anonymous):

ok what about 67.25?

OpenStudy (anonymous):

i put the numbers in incorrectly try this one http://www.wolframalpha.com/input/?i=arccos%28%2828%5E2%2B15%5E2-34%5E2%29%2F%282*15*28%29%29

OpenStudy (anonymous):

looks like 100.1 degrees from wolfram

OpenStudy (anonymous):

ok sweet would 100.08 work?

OpenStudy (anonymous):

according to wolfram it is 101.1

OpenStudy (anonymous):

ok ok its just a matter of how you round it:)

OpenStudy (anonymous):

here is a picture of the triangle http://www.wolframalpha.com/input/?i=triangle+28%2C+15%2C+34+angles+